### v1

Abstract
Spectrally truncated three-dimensional (3D) Euler flows reach an absolute equilibrium. Yet they show non-trivial temporal correlations, similar to the ones of turbulent flows. In this work, we investigate Navier-Stokes equation in equilibrium conditions, both at finite viscosity and in the inviscid limit. We determine the space-time velocity correlations, using functional renormalisation group and direct numerical simulations. We characterise the crossover in the dynamical scaling from $\tau\sim 1/(\nu k^{2})$ at finite viscosity to $\tau\sim 1/(u_{\rm rms}k)$ in the inviscid limit, where $\tau$ is the decorrelation time. We show that, although the static properties of spectrally truncated Euler flows radically differ from their hydrodynamical limit, they share a similar emergent $k^{-1}$ dynamical scaling.
1 Introduction
Since the pioneering works of Kolmogorov in 1941, deriving the statistical theory of turbulence from the fundamental hydrodynamical equations has remained an unsolved challenge. To gain some understanding, simple models, such as the three-dimensional (3D) Euler equation or the one-dimensional (1D) Burgers equation, have played a pivotal role. A noticeable feature of these models is that, while the continuous equations may generate singularities in finite time [1, 2, 3], their spectrally (Galerkin)-truncated versions lead to thermalisation to an absolute equilibrium state [4, 5, 6]. The static properties of the continuous and spectrally-truncated models thus radically differ. For the 3D Euler flow, the continuous equation is expected to yield a turbulent non-equilibrium state with Kolmogorov kinetic energy spectrum $E(k)\sim k^{-5/3}$, whereas the spectrally-truncated equation leads to the equilibrium state characterised by an equipartition energy spectrum $E(k)\sim k^{d-1}$. Note that the equipartition spectrum can also be observed in turbulent Navier-Stokes flows in the presence of thermal noise at sufficiently small scales in the dissipative range [7].
The thermalisation process to absolute equilibrium has been thoroughly characterised, and was shown to involve localised oscillatory structures called tygers, both for the 1D Burgers and for the 3D Euler spectrally-truncated equations [8, 9, 10, 11]. The emergence of these tygers deeply imprints the dynamics, yielding a relaxation time $\tau(k)\sim k^{-z}$ with $z=1$, both in the 1D Burgers [8] and in the 3D Euler [12, 13] equations. Interestingly, when a thermal noise is added to the deterministic 1D Burgers equation, the latter maps to the Kardar-Parisi-Zhang (KPZ) equation [14], a celebrated Langevin-type model in statistical physics which features non-equilibrium scaling phenomena [15, 16]. It was shown in direct numerical simulations that the stochastic 1D Burgers–KPZ equation also exhibits a decorrelation time scaling $\tau(k)\sim k^{-1}$ in the inviscid limit, as its deterministic counterpart [17, 18].
This finding weaves a deep connection between statistical physics and deterministic hydrodynamics. It turns out that this link has not been explored yet for the spectrally-truncated 3D Euler equation. The stochastic version of the 3D Navier-Stokes equation in the presence of thermal noise was introduced in the seminal paper by Forster, Nelson, Stephen (FNS) [19, 20], and named model A. The authors showed that this equation also leads to relaxation to an absolute equilibrium state, and thus model A corresponds to a Langevin model for a fluid near thermal equilibrium. Let us notice that such studies of near-equilibrium (compressible) fluids date back to the early works by Landau and Placzek [21, 22] and have important applications. However, the stochastic Euler dynamics was not discussed by FNS. In this paper, we provide the missing link, by investigating the inviscid limit of model A, using both numerical simulations and functional renormalisation group (FRG).
For the numerical part, we perform direct numerical simulations of model A of FNS, ie we integrate the spectrally-truncated stochastic 3D Navier-Stokes equation, in a Taylor-Green configuration, which allows us to access large system sizes. We compute the spatio-temporal two-point correlation function for different microscopic viscosities to determine the behaviour of the half decay time $\tau_{1/2}(k)$. We find that it also displays the $z=1$ scaling in the inviscid limit, as for the deterministic 3D Euler dynamics.
To present the FRG part, let us first summarise previous achievements on which this work builds. For the stochastic 1D Burgers–KPZ equation, the origin of the $z=1$ scaling has been recently elucidated using FRG [23]. Indeed, it was shown to be controlled by a new fixed point of the Burgers–KPZ equation termed inviscid Burgers (IB) fixed point, and characterised by a critical dynamical exponent $z=1$, differing from the well-known KPZ value $z=3/2$. The FRG is a field-theoretical method based on the continuous equations. This salient scaling of the decorrelation time is not affected by the spectral truncation, although the static properties are drastically altered. Let us notice that for the 3D Navier-Stokes equation in the turbulent state, the emergence of a $z=1$ scaling is well-known phenomenologically, and interpreted as a consequence of random sweeping effect [24, 25, 26, 27, 28, 29, 30, 31]. For the forced Navier-Stokes, a rigorous derivation of this $z=1$ was provided using FRG [33, 34, 35].
In this work, we use FRG to revisit model A of FNS, focusing in particular on its inviscid limit. We show that, beside the attractive (IR) fixed point identified by FNS, there exists another repulsive (UV) fixed point, associated with zero viscosity, which we call inviscid FNS (IFNS), in analogy with the IB fixed point for the stochastic Burgers–KPZ equation, and which is also characterised by a $z=1$ critical dynamical exponent. Thus, for both the stochastic 1D Burgers and 3D Navier-Stokes equations in equilibrium conditions, a non-trivial behaviour emerges in the inviscid limit, which is distinct from the KPZ or FNS one. Moreover, we show that this behaviour is also relevant in the viscous case, where the corresponding scaling is observable at large momentum, over a range that extends when the viscosity is further reduced.
The remainder of the paper is organised as follows. In Sec. 2, we introduce model A of FNS, and review the main results obtained by FNS relevant for this work. In Sec. 3, we briefly present the FRG formalism, and derive the flow equations for model A. We establish the fixed point structure of model A, and show that the dynamical exponent at the IFNS fixed point is $z=1$, both through a numerical integration of the complete flow equations within some approximation, and from an analytic solution of exact asymptotic flow equations in the large momentum limit. In Sec. 4, we present the results of the direct numerical simulations of model A, and compare with the FRG results. Finally, our conclusion is given in Sec. 5.
2 Stochastic Navier-Stokes equation
2.1 Model A of FNS
We consider the stochastic Navier-Stokes equation for an incompressible fluid in dimension $d$
$\partial_{t}\mathbf{v}+\lambda(\mathbf{v}\cdot\mathbf{\nabla})\mathbf{v}$ $=-\dfrac{1}{\rho}\mathbf{\nabla}\pi+\nu\nabla^{2}\mathbf{v}+\mathbf{f}\,$ (1)
$\mathbf{\nabla}\cdot\mathbf{v}$ $=0\,,$ (2)
where $\mathbf{v}(t,\mathbf{x})$ is the fluid velocity field, $\pi(t,\mathbf{x})$ the pressure field, $\nu$ the (microscopic) kinematic viscosity of the fluid, $\rho$ its density, and $\lambda$ a real parameter introduced for convenience and eventually set to unity. The field $\mathbf{f}$ is a stochastic forcing with Gaussian distribution of zero mean and of two-point correlations given in Fourier space by
$\big{\langle}f_{\alpha}(\omega,\mathbf{q})f_{\beta}(\omega^{\prime},\mathbf{q}^{\prime})\big{\rangle}=2\bar{D}(|\mathbf{q}|)P^{\perp}_{\alpha\beta}(\mathbf{q})(2\pi)^{d+1}\delta(\omega+\omega^{\prime})\delta^{d}(\mathbf{q}+\mathbf{q}^{\prime})\,,$ (3)
where $P^{\perp}_{\alpha\beta}(\mathbf{q})=\delta_{\alpha\beta}-\frac{q_{\alpha}q_{\beta}}{q^{2}}$ is the transverse (Leray) projector. The Fourier conventions used throughout this paper are
$g(\omega,\mathbf{q})$ $={\cal F}\big{[}g(t,\mathbf{x})\big{]}=\int_{t,\mathbf{x}}\,g(t,\mathbf{x})\,e^{i(\omega t-\mathbf{p}\cdot\mathbf{x})}\,,$
$g(t,\mathbf{x})$ $={\cal F}^{-1}\big{[}g(\omega,\mathbf{q})\big{]}=\int_{\omega,\mathbf{q}}\,g(\omega,\mathbf{q})\,e^{-i(\omega t-\mathbf{p}\cdot\mathbf{x})}\,,$ (4)
with the short-hand notations
$\int_{t,\mathbf{x}}\equiv\int\text{d}^{d}\mathbf{x}\text{d}t\,,\quad\int_{\omega,\mathbf{q}}\equiv\int\dfrac{\text{d}^{d}\mathbf{q}}{(2\pi)^{d}}\dfrac{\text{d}\omega}{2\pi}\,.$ (5)
Because of translational invariance in space and time, the total frequency and momentum are conserved. Thus, the Fourier transform of a generic $n$-point function depends on only $n-1$ frequencies and momenta since the last ones are fixed by conservation. We denote with an overbar the Fourier transform without the factor $(2\pi)^{d+1}\delta\left(\sum_{k=1}^{n}\omega_{k}\right)\delta^{d}\left(\sum_{k=1}^{n}\mathbf{q}_{k}\right)$ stemming from the conservation laws.
The choice $\bar{D}(q)=Dq^{2}$ models a thermal noise, and the set of equations (1)–(3) identifies in this case with model A of FNS [19, 20]. It corresponds to a Langevin dynamics for a fluid near thermal equilibrium. In particular, it satisfies a fluctuation-dissipation theorem which fixes the amplitude of the noise to $D=\frac{\nu}{\rho}k_{B}T$ where $T$ is the temperature of the fluid and $k_{B}$ the Boltzmann constant. This system relaxes at long times to an absolute equilibrium state, with equipartition of energy, yielding a kinetic energy spectrum $\bar{E}(k)\sim k^{d-1}$. In this state, the velocity has a Gaussian distribution and its two-point correlation function takes the scaling form
$\bar{G}^{(2,0)}_{\alpha\beta}(\omega,\mathbf{q})\equiv{\cal F}\big{[}\big{\langle}v_{\alpha}(t,\mathbf{x})v_{\beta}(t^{\prime},\mathbf{x}^{\prime})\big{\rangle}_{c}\big{]}=P^{\perp}_{\alpha\beta}(\mathbf{q})q^{-2}\Phi(\omega/q^{z})$ (6)
where the subscript $c$ stands for connected, and where $z$ is the dynamical critical exponent. For model A, one simply has $z=2$, characteristic of a diffusive behaviour, with the universal scaling function $\Phi$ given by
$\Phi(x)=\chi\nu_{\rm eff}/(x^{2}+\nu_{\rm eff}^{2})\,,\qquad\hbox{with}\qquad\chi=D/\nu=k_{B}T/\rho\,,$ (7)
and denoting $\nu_{\rm eff}$ the effective (measured) viscosity [20].
The relaxation process towards the equilibrium state gives rise to long-time tails, a phenomenon early discussed in the litterature [36, 20]. Casting the correlation function in the form
$\bar{G}^{(2,0)}_{\alpha\beta}(\omega,\mathbf{q})=2P^{\perp}_{\alpha\beta}(\mathbf{q})\Re\left[\dfrac{\chi}{-i\omega+q^{2}\nu_{R}(\omega,q)}\right]$ (8)
to define the renormalised viscosity $\nu_{R}(\omega,q)$, the long-time tail corresponds to a slow decay as $\nu_{R}(t,q=0)\sim t^{-d/2}$ for $d>2$, with logarithmic corrections in $d=2$.
The constitutive equations of model A can be rescaled, such that there remains a single parameter $\bar{\lambda}=\lambda D^{1/2}/{\nu}^{3/2}$. As emphasised by FNS [20], the absolute equilibrium state corresponds to a vanishing renormalised $\bar{\lambda}$. This is why the correlations (6)–(7) are similar to results from linearised hydrodynamics. Moreover, the near equilibrium dynamics can be captured using perturbation theory, which is organised as a power series in $\bar{\lambda}$. However, this approach is not suitable to study fully developed turbulence, which corresponds to a non-trivial steady-state with a finite $\bar{\lambda}$ and non-Gaussian fluctuations. The perturbation theory in $\bar{\lambda}$ obviously also fails to study the inviscid limit $\nu\to 0$, even in the equilibrium state. For this, one has to resort to non-pertubative methods, such as the FRG which we explain in Sec. 3.
2.2 Action and symmetries for model A
In order to apply renormalisation group techniques, one can cast the stochastic equation into a path integral formulation, following the Martin-Siggia-Rose-Janssen-de Dominicis formalism [37, 38, 39]. One obtains from the stochastic Navier-Stokes equation with thermal noise (1)–(3) the following generating functional and action [35]
${\cal Z}[\mathbf{j},\bar{\mathbf{j}},K,\bar{K}]$ $=\int{\cal D}\mathbf{v}{\cal D}\bar{\mathbf{v}}{\cal D}\pi{\cal D}\bar{\pi}\,e^{-{\cal S}[\mathbf{v},\bar{\mathbf{v}},\pi,\bar{\pi}]+\int_{t,\mathbf{x}}\big{\{}\mathbf{j}\cdot\mathbf{v}+\bar{\mathbf{j}}\cdot\bar{\mathbf{v}}+K\pi+\bar{K}\bar{\pi}\big{\}}}$ (9)
${\cal S}[\mathbf{v},\bar{\mathbf{v}},\pi,\bar{\pi}]$ $=\int_{t,\mathbf{x}}\Bigg{\{}\bar{v}_{\alpha}\Big{(}\partial_{t}v_{\alpha}+\lambda v_{\beta}\partial_{\beta}v_{\alpha}+\dfrac{1}{\rho}\partial_{\alpha}\pi-\nu\partial^{2}v_{\alpha}\Big{)}+\bar{\pi}\partial_{\gamma}v_{\gamma}-D\big{(}\partial_{\alpha}\bar{v}_{\beta}\big{)}^{2}\Bigg{\}}\,,$ (10)
where $\bar{\mathbf{v}}$ and $\bar{\pi}$ are the response velocity and response pressure fields, and $\mathbf{j},\bar{\mathbf{j}},K,\bar{K}$ are the sources linearly coupled to the fields. The functional ${\cal W}=\ln{\cal Z}$ is the generating functional of connected correlation functions – which are the equivalent, for fluctuating fields instead of random variables, to cumulants. Indeed, the connected correlation and response functions are obtained as
$G^{(m,n)}_{\alpha_{1}...\alpha_{m+n}}(t_{1},\mathbf{x}_{1};\ldots;t_{m+n},\mathbf{x}_{m+n})\equiv\dfrac{\delta^{m+n}{\cal W}}{\delta j_{\alpha_{1}}(t_{1},\mathbf{x}_{1})\ldots\delta\bar{j}_{\alpha_{m+n}}(t_{m+n},\mathbf{x}_{m+n})}\Big{|}_{\mathbf{j},\bar{\mathbf{j}}=0}\,,$ (11)
where the $m$ first derivatives are with respect to $\mathbf{j}$ sources and the $n$ last with respect to $\bar{\mathbf{j}}$ sources.
Another important generating functional is the Legendre transform of ${\cal W}$, which is called the effective action $\Gamma$ and is defined by
$\Gamma[\Psi]=\underset{\mathcal{J}}{\sup}\left[-\mathcal{W}[\mathcal{J}]+\int_{t,\mathbf{x}}\mathcal{J}\Psi\right]\,,\qquad\hbox{where}\quad\Psi_{i}\equiv\langle\Phi_{i}\rangle_{\cal J}=\dfrac{\delta{\cal W}}{\delta{\cal J}_{i}}\,,$ (12)
with $\Phi=(\mathbf{v},\bar{\mathbf{v}},\pi,\bar{\pi})$ denoting the multiplet of fields, $\mathcal{J}=(\mathbf{j},\bar{\mathbf{j}},K,\bar{K})$ the associated sources, and $\Psi=(\mathbf{u},\bar{\mathbf{u}},\Pi,\bar{\Pi})$ the expectation values of the fields in the presence of the sources ${\cal J}$. One can obtain from $\Gamma$ the vertex functions, defined as
$\Gamma^{(m,n)}_{\alpha_{1}...\alpha_{m+n}}(t_{1},\mathbf{x}_{1};\ldots;t_{m+n},\mathbf{x}_{m+n})\equiv\dfrac{\delta^{m+n}\Gamma}{\delta u_{\alpha_{1}}(t_{1},\mathbf{x}_{1})\ldots\delta\bar{u}_{\alpha_{m+n}}(t_{m+n},\mathbf{x}_{m+n})}\Big{|}_{\mathbf{u},\bar{\mathbf{u}}=0}\,.$ (13)
The knowledge of the set of $G^{(m,n)}$, or equivalently of the set of $\Gamma^{(m,n)}$, provides the statistical theory for the model under study.
Let us now briefly explain how the symmetries lead to exact relations between these functions, which are called Ward identities, and which can be very efficiently exploited within the FRG formalism. In fact, it turns out to be particularly useful to consider extended symmetries, which are continuous infinitesimal transformations which do not leave the action strictly invariant, but yields a variation at most linear in the fields. One can indeed derive from such extended symmetries exact Ward identities, which deeply constrain the set of $G^{(m,n)}$ (or equivalently $\Gamma^{(m,n)}$[40].
This procedure has been expounded in details, in eg. Refs. [34, 35] for the Navier-Stokes equation, or in eg. Refs. [41, 23] for the related Burgers–KPZ equation. Hence, we only provide here a summary, and refer the interested reader to these references for technical details. The first noticeable extended symmetries are shifts of the pressure fields, $\pi(t,\mathbf{x})\to\pi(t,\mathbf{x})+\epsilon(t,\mathbf{x})$, and $\bar{\pi}(t,\mathbf{x})\to\bar{\pi}(t,\mathbf{x})+\bar{\epsilon}(t,\mathbf{x})$, with $\epsilon(t,\mathbf{x})$ and $\bar{\epsilon}(t,\mathbf{x})$ infinitesimal scalar functions of time and coordinate. The corresponding Ward identities impose that the variation of the effective action $\Gamma$ under these transformations is equal to that of the action ${\cal S}$, which entails that the whole pressure sector is not renormalised. A second extended symmetry is the time-dependent Galilean symmetry, which corresponds to the transformation
$\delta v_{\alpha}(t,\mathbf{x})=-\dot{\epsilon}_{\alpha}(t)+\epsilon_{\beta}(t)\partial_{\beta}v_{\alpha}(t,\mathbf{x})\,,$ $\quad\delta\pi(t,\mathbf{x})=\epsilon_{\beta}(t)\partial_{\beta}\pi(t,\mathbf{x})\,,$
$\delta\bar{v}_{\alpha}(t,\mathbf{x})=\epsilon_{\beta}(t)\partial_{\beta}\bar{v}_{\alpha}(t,\mathbf{x})\,,$ $\quad\delta\bar{\pi}(t,\mathbf{x})=\epsilon_{\beta}(t)\partial_{\beta}\bar{\pi}(t,\mathbf{x})\,,$ (14)
with $\epsilon(t)$ an infinitesimal vectorial function of time. The corresponding Ward identities exactly relate a vertex function $\Gamma^{(m+1,n)}$ with one vanishing velocity momentum to lower order ones $\Gamma^{(m,n)}$. We here simply give the first two identities
$\bar{\Gamma}^{(2,1)}_{\alpha\alpha_{1}\alpha_{2}}(\omega,0;\omega_{1},\mathbf{p}_{1})=-\frac{p_{1\alpha}}{\omega}\left[\bar{\Gamma}^{(1,1)}_{\alpha_{1}\alpha_{2}}(\omega_{1}+\omega,\mathbf{p}_{1})-\bar{\Gamma}^{(1,1)}_{\alpha_{1}\alpha_{2}}(\omega_{1},\mathbf{p}_{1})\right],$
$\bar{\Gamma}^{(1,2)}_{\alpha\alpha_{1}\alpha_{2}}(\omega,0;\omega_{1},\mathbf{p}_{1})=-\frac{p_{1\alpha}}{\omega}\left[\bar{\Gamma}^{(0,2)}_{\alpha_{1}\alpha_{2}}(\omega_{1}+\omega,\mathbf{p}_{1})-\bar{\Gamma}^{(0,2)}_{\alpha_{1}\alpha_{2}}(\omega_{1},\mathbf{p}_{1})\right]\,,$ (15)
general expressions for arbitrary $(m,n)$ can be found in Ref. [35].
The third important extended symmetry is the time-dependent response shift, corresponding to the transformation
$\delta\bar{v}_{\alpha}(t,\mathbf{x})=\bar{\epsilon}_{\alpha}(t)\,,\quad\delta\bar{\pi}(t,\mathbf{x})=v_{\beta}(t,\mathbf{x})\bar{\epsilon}_{\beta}(t)\,,$ (16)
with $\bar{\epsilon}(t)$ an infinitesimal vectorial function of time. The corresponding Ward identities exactly impose that a vertex function $\Gamma^{(m,n)}$ with one vanishing response velocity momentum vanish, except the two lowest order ones, which are fixed to
$\bar{\Gamma}_{\alpha\beta}^{(1,1)}(\omega,\mathbf{p}=\mathbf{0})=i\omega\delta_{\alpha\beta},$
$\bar{\Gamma}_{\alpha\beta\gamma}^{(2,1)}(\omega_{1},\mathbf{p}_{1},\omega_{2},-\mathbf{p}_{1})=ip_{1}^{\alpha}\delta_{\beta\gamma}-ip_{1}^{\beta}\delta_{\alpha\gamma}\,.$ (17)
Finally, model A possesses an additional symmetry compared to Navier-Stokes equation (without thermal noise), which is a time-reversal symmetry, corresponding to the discrete transformation
$v_{\alpha}(t,\mathbf{x})\to-v_{\alpha}(-t,\mathbf{x})\,,\quad\pi(t,\mathbf{x})\to\pi(-t,\mathbf{x})\,,$
$\bar{v}_{\alpha}(t,\mathbf{x})\to\bar{v}_{\alpha}(-t,\mathbf{x})-\dfrac{\nu}{D}v_{\alpha}(-t,\mathbf{x})\,,\quad\bar{\pi}(t,\mathbf{x})\to-\bar{\pi}(-t,\mathbf{x})\,.$ (18)
This symmetry also yield constraints on the vertex functions, in particular
$2\Re{\rm e}\Big{[}\bar{\Gamma}^{(1,1)}_{\alpha\beta}(\omega,\mathbf{p})\Big{]}=-\dfrac{\nu}{D}\bar{\Gamma}^{(0,2)}_{\alpha\beta}(\omega,\mathbf{p})\,,$ (19)
which is a fluctuation-dissipation relation. All these identities are exploited in the following to devise suitable approximation schemes in the FRG formalism.
3 Functional renormalisation group
3.1 FRG formalism
We briefly lay out the basis of the FRG formalism, referring the reader to comprehensive introductions, such as [42, 43, 44, 45], for more details. The FRG is based on the idea of Wilsonian RG. Instead of averaging over the fluctuating fields $\Phi$ in the path integral (9) at all (momentum) scales simultaneously, one only integrates up to a certain RG scale $\kappa$. One then considers how this modified theory evolves under an infinitesimal change of the RG scale, which gives rise to a deterministic differential equation, where the role ot time is played by the RG scale, and the corresponding trajectories are in an abstract theory space, and connect a given microscopic model to its effective theory at large distance and time. From the technical point of view, this is achieved in FRG through the introduction of a scale-dependent ‘weight’ $\exp(-\Delta\mathcal{S}_{\kappa})$ in the functional integral to suppress fluctuations below the RG scale $\kappa$:
$\mathcal{Z}_{\kappa}=\int{\mathrm{D}}\Phi\exp\left(-\mathcal{S}[\Phi]-\Delta\mathcal{S}_{\kappa}[\Phi]+\int_{t,\mathbf{x}}\mathcal{J}\Phi\right),$ (20)
where $\Delta\mathcal{S}_{\kappa}[\Phi]\equiv\frac{1}{2}\int_{t,\mathbf{x},\mathbf{x}^{\prime}}\Phi(t,\mathbf{x})\mathcal{R}_{\kappa}(|\mathbf{x}-\mathbf{x}^{\prime}|)\Phi(t,\mathbf{x}^{\prime})$. The kernel $\mathcal{R}_{\kappa}$ is called ‘regulator’ and must satisfy the following properties (in Fourier space): (i) $\mathcal{R}_{\kappa}(\mathbf{p})\stackrel{{\scriptstyle\kappa\rightarrow\Lambda}}{{\sim}}\Lambda^{2}$ such that all fluctuations are frozen at the initial (microscopic) RG scale $\Lambda$; (ii) $\mathcal{R}_{\kappa}(\mathbf{p})\stackrel{{\scriptstyle\kappa\rightarrow 0}}{{\longrightarrow}}0$ such that all fluctuations are averaged over at the final (large-distance) RG scale; (iii) $\mathcal{R}_{\kappa}(\mathbf{p})\stackrel{{\scriptstyle|\mathbf{p}|\ll\kappa}}{{\sim}}\kappa^{2}$ such that at a scale $\kappa$, the contribution of the ‘slow’ modes is suppressed; (iv) $\mathcal{R}_{\kappa}(\mathbf{p})\stackrel{{\scriptstyle|\mathbf{p}|\gg\kappa}}{{\longrightarrow}}0$ such that at a scale $\kappa$, the ‘fast’ modes are unaltered and averaged over. In the presence of the regulator, the definition (12) is slightly modified to
$\Gamma_{\kappa}[\Psi]+\Delta\mathcal{S}_{\kappa}[\Psi]=\underset{\mathcal{J}}{\sup}\left[-\mathcal{W}_{\kappa}[\mathcal{J}]+\int_{t,\mathbf{x}}\mathcal{J}\Psi\right]\,,$ (21)
where ${\cal W}_{\kappa}=\ln{\cal Z}_{\kappa}$, and $\Gamma_{\kappa}$ is called the effective average action. In the large-$\kappa$ limit, the microscopic action $\Gamma_{\Lambda}\equiv\mathcal{S}$ is recovered since all the fluctuations are frozen, while in the small-$\kappa$ limit, the full effective action $\Gamma_{0}\equiv\Gamma$, encompassing all the fluctuations at all scales, is obtained [44, 45]. This provides $\Gamma_{\kappa}$ with a physical interpretation as the effective theory at scale $\kappa$, interpolating from the microscopic theory at $\kappa=\Lambda$ to the complete effective theory when $\kappa\to 0$. The very idea of the Wilsonian RG is thus realised through this formalism in a very intuitive way. The evolution of $\Gamma_{\kappa}$ with the RG scale in between these two limits is given by the exact Wetterich equation [46, 47, 48]
$\partial_{\kappa}\Gamma_{\kappa}=\frac{1}{2}\textrm{tr}\,\int_{\omega,\mathbf{q}}\partial_{\kappa}\mathcal{R}_{\kappa}\,\mathcal{G}_{\kappa}\,,\qquad\hbox{with}\quad\mathcal{G}_{\kappa}\equiv\left(\Gamma_{\kappa}^{(2)}+\mathcal{R}_{\kappa}\right)^{-1}\,,$ (22)
where the trace means summation over all fields, $\Gamma_{\kappa}^{(2)}$ is the Hessian of $\Gamma$, and $\mathcal{G}_{\kappa}$ the propagator matrix.
The flow equation (22) is exact, but as it is a functional partial differential equation, it cannot be solved exactly in general. In the following, we present two complementary approaches to treat these equations for model A and investigate its inviscid limit. One of them consists in devising an Ansatz for $\Gamma_{\kappa}$, which allows one to project the flow onto a set of chosen renormalised functions, which are then numerically solved, as done in eg. Ref. [49]. This approach is particularly accurate to describe the small-momentum behaviour. It allows us to first qualitatively establish the fixed point structure of model A, evidencing the existence of the IFNS fixed point. The other approach consists in expanding the flow equation for the correlation function at large momentum. The leading order of this equation can then be closed exactly using the Ward identities stemming from the extended symmetries. Moreover, it can be solved analytically, yielding the exact asymptotic expression of the correlation function at large momentum. This approach is appropriate to describe the large-momentum behaviour. It allows us to show that $z=1$ exactly at the IFNS fixed point. Finally, to obtain a quantitative picture over the whole range of momenta, we consider an advanced numerical method which allows us to solve simultaneously the two sets (large- and small-momentum) of flow equations and compute the full correlation function and the associated half decay frequency $\omega_{1/2}(p)$. We start by presenting the large-momentum equations to establish the general result $z=1$, before turning to the small-momentum one to show the existence of the new IFNS fixed point.
3.2 Exact closure at large momentum
The exact FRG flow equations for any vertex function $\Gamma_{\kappa}^{(m,n)}$ can be obtained by taking functional derivatives of Eq. (22). It turns out that it can be simplified in a controlled way in the limit of large momenta. This originates in the very structure of these flow equations, in particular in the presence of the regulator term ${\partial}_{s}\mathcal{R}_{\kappa}(\mathbf{q})$ in the loop integral. It implies that the integral over the internal momentum $\mathbf{q}$ is cut to values $|\mathbf{q}|\lesssim\kappa$. Thus, in the limit where all the external momenta are large $|\mathbf{p}_{i}|\gg\kappa$ for all $i$, one has $|\mathbf{q}|\ll|\mathbf{p}_{i}|$. Because analyticity of the vertices is guaranteed at all scales $\kappa$ [45], the vertices appearing in the integrand can be Taylor expanded in powers of $\mathbf{q}$, since $|\mathbf{q}|$ is always compared to one of the $|\mathbf{p}_{i}|$ and $|\mathbf{q}|/|\mathbf{p}_{i}|\to 0$ for all $i$. This expansion becomes exact in the limit where all the $\mathbf{p}_{i}$ (and their partial sums) are infinite and in the vicinity of a fixed point [50, 51, 52].
While the former property always holds, an exceptional feature occurs in the context of turbulence, which stems from the existence of the extended symmetries highlighted in Sec. 22.2. Indeed, the set of related Ward identities (15), (17) (and similar higher-order ones) appears to fix all the vertices with one zero momentum in terms of lower-order vertices. These are precisely the configurations which arise in the flow equations expanded at large momenta ($|\mathbf{q}|/|\mathbf{p}_{i}|\to 0$), which can thus be closed exactly using the Ward identities. Moreover, the resulting closed flow equations can be solved exactly in the vicinity of a fixed point. This exact closure was first achieved for the forced (non-equilibrium) Navier-Stokes equation in three, and then two dimensions, and for passively advected scalars [49, 33, 34, 53, 54]. It was also demonstrated for the one-dimensional, and $d$-dimensional, Burgers–KPZ equation [23, 55]. We refer the readers to these references or to the review [35] for the complete proof.
The crucial point is that the exact FRG flow equations for the $n$-point correlation functions are very generic, as well as their large momentum expansion. Moreover, their exact closure only relies on the extended symmetries, and not on the details of the specific hydrodynamical equation, in particular not on the precise form of the noise or forcing term provided it has a Gaussian distribution, since this term does not enter the Ward identities (15), (17). In this respect, the large momentum equations of model A are the same as the ones obtained for the forced Navier-Stokes, except that there is only one independent two-point function because of the identity (19) stemming from time-reversal symmetry. We give it for the transverse correlation function $\bar{C}(\varpi,\mathbf{p})\equiv P_{\alpha\beta}^{\perp}(\mathbf{p})\bar{G}_{\alpha\beta}^{(2,0)}(\varpi,\mathbf{p})$ (the longitudinal part vanishes because of incompressibility) [34, 35]
$\partial_{s}\bar{C}_{\kappa}(\varpi,\mathbf{p})=\dfrac{d-1}{d}{p^{2}}\int_{\omega}\dfrac{\bar{C}_{\kappa}(\omega+\varpi,\mathbf{p})-\bar{C}_{\kappa}(\varpi,\mathbf{p})}{\omega^{2}}\,{J}_{\kappa}(\omega)\,,\qquad{J}_{\kappa}(\omega)\equiv\tilde{\partial}_{s}\int_{\mathbf{q}}\bar{C}_{\kappa}(\omega,\mathbf{q})\,,$ (23)
with $\partial_{s}\equiv\kappa\partial_{\kappa}$, and where ~$\tilde{\partial}_{s}$ only acts on the scale dependence of the regulator. This equation can be solved at a fixed point. Fourier transforming back to time delay, the solution reads [34, 35]
$\bar{C}(t,\mathbf{p})=\bar{C}(0,\mathbf{p}){\times}\begin{cases}\exp\left(-\mu_{0}(pt)^{2}\right),\quad t\ll\tau_{c}\\ \exp\left(-\mu_{\infty}p^{2}|t|\right),\quad t\gg\tau_{c}\end{cases}$ (24)
where $\mu_{0}$ and $\mu_{\infty}$ are non-universal constants (related to the fixed-point values of $J_{\kappa}(0)$ and $\int_{\omega}J_{\kappa}(\omega)$ respectively) and $\tau_{c}$ is a characteristic time scale. This expression shows that the correlation function takes at small time delays a scaling form similar to (6) with the scaling variable $x=pt$, and thus $z=1$, different from the FNS value $z=2$. Hence, even in the viscous case where the small momenta (large distance) converge to the scaling form (6) with $z=2$, the large momenta exhibit a different scaling when the viscosity is small enough. Moreover, the solution (24) predicts a crossover at large time delays to yet another scaling regime, with an exponent $z=2$, but which differs from the viscous regime since $\mu_{\infty}$ does not depend on $\nu$, and it emerges at large momentum (not at large distances). Let us note that the solution (24) is an exact asymptotic result, but it does not give quantitative information about when (i.e. for how large momentum, and how small viscosity) it holds. This issue will be investigated in Sec. 33.4.
As previously mentioned, the result (24) is very general, it also holds for the large-momentum behaviour of the forced Navier-Stokes equation or passive scalar turbulence. It has been investigated in these contexts through direct numerical simulations. While the small-time $z=1$ decay is easily observed, eg. [28, 31], the large-time $z=2$ decay has remained elusive so far. Indeed, this regime turns out to be hard to resolve in numerical simulations since the initial Gaussian decay is very fast and the large-time correlations fall below the noise level. This regime has been evidenced so far only for a generalisation of Kraichnan’s model for passive scalars in Ref. [32]. We show in Sec. 4 that for model A, we manage to evidence both the small and the large time regimes in our direct numerical simulations.
3.3 FRG flow equations for model A
The asymptotic expression (24) holds for large $p$. However, the constants $\mu_{0}$, $\mu_{\infty}$ are related to the fixed-point values of $J_{\kappa}(\omega)$ in (23[34]. They hence primarily depend on the small $\mathbf{q}$ region, and their calculation requires the integration of the whole flow. An approximation which is accurate to describe the small momentum sector is to devise an ansatz for $\Gamma_{\kappa}$. Taking into account all the symmetries and associated non-renormalisation theorems, the appropriate ansatz reads [49]
$\Gamma_{\kappa}=\int_{t,\mathbf{x}}\Bigg{\{}\bar{v}_{\alpha}\Big{(}\partial_{t}v_{\alpha}+\lambda v_{\beta}\partial_{\beta}v_{\alpha}+\frac{1}{\rho}\partial_{\alpha}\pi-\nu_{\kappa}(\partial_{t},\mathbf{\nabla})\partial^{2}v_{\alpha}\Big{)}+\bar{\pi}\partial_{\gamma}v_{\gamma}-\chi\nu_{\kappa}(\partial_{t},\mathbf{\nabla})\big{(}\partial_{\alpha}\bar{v}_{\beta}\big{)}^{2}\Bigg{\}}\,\,,$ (25)
where in Fourier space $\bar{\nu}_{\kappa}(\omega,\mathbf{p})$ is a whole function of $\omega$ and $\mathbf{p}$, which identifies with the renormalised viscosity defined by (8). However, in contrast with perturbative calculations, the FRG allows one to compute the flow of this whole function (hence the F in FRG), and not only renormalised parameters as standard dynamical RG. The ansatz (25) encompasses the most general two-point functions, while it neglects all higher-order vertices but $\Gamma_{\kappa}^{(2,1)}$ which is present at the bare level and is not renormalised. It was shown to yield accurate results for 3D and 2D turbulence [35].
Projecting the exact Wetterich equation (22) onto the ansatz (25), one obtains the flow equation of $\bar{\nu}_{\kappa}(\omega,\mathbf{p})$ as
$\partial_{s}\bar{\nu}_{\kappa}(\omega,\mathbf{p})$ $=\frac{2\lambda^{2}}{(d-1)\mathbf{p}^{2}}\int_{\varpi,\mathbf{q}}\Bigg{\{}\partial_{s}R_{\kappa}(\mathbf{q})\dfrac{\varpi^{2}-\tilde{\nu}_{\kappa}(\varpi,\mathbf{q})^{2}}{\big{(}\varpi^{2}+(\tilde{\nu}_{\kappa}(\varpi,\mathbf{q}))^{2}\big{)}^{2}}\dfrac{\tilde{\nu}_{\kappa}(\omega+\varpi,\mathbf{p}+\mathbf{q})}{(\omega+\varpi)^{2}+(\tilde{\nu}_{\kappa}(\omega+\varpi,\mathbf{p}+\mathbf{q}))^{2}}$
×$\times\Big{[}\Big{(}2\mathbf{p}^{2}-\frac{(\mathbf{p}\cdot(\mathbf{p}+\mathbf{q}))^{2}}{(\mathbf{p}+\mathbf{q})^{2}}-\frac{(\mathbf{p}\cdot\mathbf{q})^{2}}{\mathbf{q}^{2}}\Big{)}(d-1)$
$+2\frac{1}{\mathbf{q}^{2}(\mathbf{p}+\mathbf{q})^{2}}\Big{(}\mathbf{q}^{2}-\frac{(\mathbf{p}\cdot\mathbf{q})^{2}}{\mathbf{p}^{2}}\Big{)}\Big{(}\mathbf{p}^{2}\mathbf{p}\cdot\mathbf{q}+2(\mathbf{p}\cdot\mathbf{q})^{2}-\mathbf{p}^{2}\mathbf{q}^{2}\Big{)}\Big{]}\Bigg{\}}\,,$ (26)
where $R_{\kappa}$ denotes the component of the regulator matrix ${\cal R}_{\kappa}$, $R_{\kappa}\equiv[{\cal R}_{\kappa}]_{12}=[{\cal R}_{\kappa}]_{21}=-[{\cal R}_{\kappa}]_{22}/2$. with $\tilde{\nu}_{\kappa}(\varpi,\mathbf{q})=\mathbf{q}^{2}\bar{\nu}_{\kappa}(\varpi,\mathbf{q})+R_{\kappa}(\mathbf{q})$. The derivation of this equation can be found in Ref. [49], simply imposing at the end the constraints ensuing from time-reversal symmetry  [^1] The flow equation (26) corresponds to Eq. (B10) of Ref. [49], setting $f_{\perp}^{D}(p)=\chi f_{\perp}^{\nu}(\mathbf{p})$ and $N_{\kappa}=\chi R_{\kappa}$ to preserve time-reversal symmetry, and restoring the frequency arguments $f_{\perp}^{\nu}(\mathbf{q})\to\tilde{\nu}_{\kappa}(\varpi,\mathbf{q})$, $f_{\perp}^{\nu}(\mathbf{p}+\mathbf{q})\to\tilde{\nu}_{\kappa}(\omega+\varpi,\mathbf{p}+\mathbf{q})$. Note that a sign typo in the last line of (B10) has been corrected.. Note that a further simplification is used for the numerical resolution of this equation, which consists in neglecting the frequency dependence of $\tilde{\nu}_{\kappa}$ within the integral, which amounts to the replacement $\tilde{\nu}(\varpi,\mathbf{q})\simeq\tilde{\nu}_{\kappa}(0,\mathbf{q})$ in the integrand.
3.4 Results from FRG
We first elucidate the fixed point structure of the FNS equation within an even simpler approximation, which consists in considering only the (non-perturbative) flow of the non-linear dimensionless coupling parameter $\bar{g}_{\kappa}=\kappa^{d-4}\chi\lambda/\nu_{\kappa}^{2}$, with $\nu_{\kappa}\equiv\bar{\nu}_{\kappa}(0,0)$. Since we are interested in the inviscid fixed point, corresponding to $\bar{g}_{\kappa}\to\infty$, we define $\hat{w}_{\kappa}=\bar{g}_{\kappa}/(1+\bar{g}_{\kappa})$. Its flow is given by
$\partial_{s}\hat{w}_{\kappa}=\hat{w}_{\kappa}(1-\hat{w}_{\kappa})\big{(}d-2+2\eta_{\kappa}\big{)}\,,\quad\hbox{with}\;\;\eta_{\kappa}=-\partial_{s}\ln\nu_{\kappa}\,.$ (27)
This equation can be straightforwardly integrated. The result in shown in Fig. 1, where it is compared with the equivalent flow for the 1D Burgers equation from Ref. [23]. From any initial condition $\hat{w}_{\Lambda}$ at $\kappa=\Lambda$ ($s=\ln(\kappa/\Lambda)=0$), the flow reaches, as $\kappa\to 0$ ($s\to-\infty$), an IR fixed point, with $\hat{w}^{\rm FNS}_{*}=0$ which is the attractive FNS fixed point. To evidence UV fixed points, one can run the flow backwards, from a very small initial $\kappa$ (very large negative initial $s$) to $\kappa\to\Lambda$ ($s\to 0$), such that the stability of the fixed points are exchanged. The backwards flow always reaches in the UV a new fixed point, with $\hat{w}^{\rm IFNS}_{*}=1$, which corresponds to the IFNS fixed point, as illustrated on Fig. 1(b). This demonstrates the existence of this fixed point. In contrast, for the 1D Burgers flow, while the flow always leads in the IR to a non-Gaussian fixed point with a finite $10<\hat{w}^{\rm KPZ}_{*}<$ 1 which is the KPZ fixed point, it can reach in the UV two different UV fixed points depending on the initial value of the coupling: the IB fixed point for $\hat{w}_{0}>\hat{w}^{\rm KPZ}_{*}$, and the EW one for $\hat{w}_{0}<\hat{w}^{\rm KPZ}_{*}$, as illustrated on Fig. 1(a).
Figure 1: FRG flows of the non-linear coupling $\hat{w}_{\kappa}$ for the stochastic (a) 1D Burgers and (b) 3D Navier-Stokes equations: in red flows from $s=\ln(\kappa/\Lambda)=0$ to $s\to-\infty$ ($\kappa\to 0$) towards the IR fixed points for different initial conditions $\hat{w}_{\Lambda}$; in blue flows from $s\gtrsim-20$ to $s\to 0$ ($\kappa\to\Lambda$) towards the UV fixed points for different initial conditions $\hat{w}_{0}$.
This unveils the whole fixed point structure of model A, which is sketched in Fig. 2, again in correspondence with the one of the Burgers–KPZ equation. The FNS fixed point describes the large distance and time, i.e. the IR modes (symbolised by the red-shaded area in the figure), of the viscous fluid at rest. The IFNS fixed point describes the inviscid fluid at equilibrium, but it also controls the large (UV) modes of a viscous fluid, the range of these modes (symbolised by the blue-shaded area in the figure) extending as the viscosity decreases.
Figure 2: Sketch of the fixed point structure of the stochastic 1D Burgers and 3D Navier-Stokes equations in the presence of thermal noise. The red points represent attractive (IR) fixed points and the blue points repulsive (UV) ones. The red-shaded (resp. blue-shaded) area hence symbolises the IR (resp. UV) modes controlled by the IR (resp. UV) fixed points.
In order to obtain more quantitative results, and in particular the correlation function, we numerically solve the complete set of flow equations (23), (26), using two complementary grids [52]. A first grid is used for the small momentum sector, where the flow equation (26) is integrated using dimensionless variables with a similar numerical scheme as described in Refs. [23, 49]. The large momentum flow equation (23) is integrated on a second grid using dimensionful variables, with the input at each scale $\kappa$ of $J_{\kappa}(\omega)$ computed on the dimensionless grid. The full numerical procedure is detailed in Ref. [56]. This scheme allows us to compute the correlation function $\bar{C}(\omega,\mathbf{p})$ on the whole momentum range. To evidence scaling regimes, we extract from it the half decay frequency defined such that $\bar{C}(\omega_{1/2}(q),\mathbf{p})=\bar{C}(0,\mathbf{p})/2$. The result is displayed on Fig. 3. Two markedly different scaling regimes are observable, thus unveiling the existence of two underlying fixed points. At small momenta, one finds $\omega_{1/2}\sim q^{z}$ with $z=2$ the FNS diffusive exponent. This region is hence controlled by the IR FNS fixed point as expected. For large viscosities (eg $\nu=1$), only this scaling is visible. For large momenta, when the viscosity decreases, one finds another scaling regime where $\omega_{1/2}\sim q^{z}$ with $z=1$. This regime is all the more extended than the viscosity is small.
Figure 3: Half decay frequency for different initial values of the viscosity. The IR modes show a diffusive FNS scaling $\omega_{1/2}\sim p^{2}$. The IFNS scaling, characterised by $\omega_{1/2}\sim p$, appears for UV modes when the viscosity decreases.
4 Results from direct numerical simulations
4.1 Numerical methods
The stochastically forced Navier-Stokes equations (1)– (3) are numerically solved using $2/3$ de-aliased Fourier spectral methods [57]. We thus use the discrete Fourier transform $\mathbf{v}(t,\mathbf{x})=\sum\hat{\mathbf{v}}(t,\mathbf{k})e^{i\mathbf{k}\cdot\mathbf{x}}$ of a 3D spatially periodic velocity field. Performing a Galerkin truncation on (1)–(3), i.e. ${\hat{\mathbf{v}}}(t,\mathbf{k})=0$ for $k_{\alpha}k_{\alpha}\geq k_{\rm max}^{2}$ (where $k_{\rm max}$ is $1/3$ of the resolution), generates the following finite system of ordinary differential equations for the complex variables ${\hat{\mathbf{v}}}(t,\mathbf{k})$ ($\mathbf{k}$ being a 3D vector of relative integers $(k_{1},k_{2},k_{3})$)
${\partial_{t}{\hat{v}}_{\alpha}(t,\mathbf{k})}=-\frac{i}{2}{\mathcal{P}}_{\alpha\beta\gamma}(\mathbf{k})\sum_{\bf p}{\hat{v}}_{\beta}(t,{\bf p}){\hat{v}}_{\gamma}(t,{\bf k-p})-\nu k_{\beta}k_{\beta}{\hat{v}}_{\alpha}(t,\mathbf{k})+i\sqrt{D}\epsilon_{\alpha\beta\gamma}k_{\beta}{\hat{f}}_{\gamma}$ (28)
where ${\mathcal{P}}_{\alpha\beta\gamma}(\mathbf{k})=k_{\beta}P^{\perp}_{\alpha\gamma}(\mathbf{k})+k_{\gamma}P^{\perp}_{\alpha\beta}(\mathbf{k})$ and the convolution in (28) is truncated (set to zero) for $k_{\alpha}k_{\alpha}\geq k_{\rm max}^{2}$.
When $D=\nu=0$, this truncated Euler equation is a time-reversible system which exactly conserves the kinetic energy $E=\sum_{k}E(t,k)$, where the energy spectrum $E(t,k)$ is defined by averaging ${\hat{\mathbf{v}}}(t,{\mathbf{k}^{\prime}})$ on spherical shells of width $\Delta k=1$,
$E(t,k)={\frac{1}{2}}\sum_{k-\Delta k/2<|{\mathbf{k}^{\prime}}|<k+\Delta k/2}|{\hat{\mathbf{v}}}({\mathbf{k}^{\prime}},t)|^{2}\,.$ (29)
At large times, the truncated 3D Euler equations is known to reach an absolute equilibrium, which is a statistically stationary exact Gaussian solution, with energy spectrum $E(k)=ck^{2}$ [12]. When $\nu$ and $D$ are non-zero, identical statistically stationary solutions also exist. A realisation of such a solution, with $u_{rms}=\sqrt{2E}=\frac{1}{2}$, is used as initial data to start our numerical simulations. When $\nu$ is non-zero, $D$ is adjusted so that the corresponding stationary solution also has $u_{rms}=\frac{1}{2}$ (see [17] for details on a similar setting for the 1D Burgers equation).
We use fields with Taylor-Green symmetries (see Ref. [13] for details). Time-stepping of the deterministic terms of (28) is performed using a $4^{\rm th}$-order Runge-Kutta method and the random forcing term is added independently. Two-dimensional slices of the 3D computed Fourier-space velocity field are output periodically every $\delta t$ to files. Post-processing the files, time-correlation functions and spatio-temporal power spectra up to the Nyquist pulsation $\omega_{\rm max}=\pi/\delta t$ are estimated using standard algorithms [58]. The runs presented here use spatial resolution of $256^{3}$ and $u_{rms}=\frac{1}{2}$.The inviscid run uses the output frequency $\delta t=0.1$ and a total integration time of $t=1024$. The viscous run with $\nu=0.016$ uses $\delta t=0.0125$ and a total integration time of $t=128$.
4.2 Scalings of correlation times
We first study the correlation function $\bar{C}(t,k)$, and the related half decay time $\tau_{1/2}(k)$. In order to visualise its scaling behaviour, we represent in Fig. 4 the map of the normalised correlations $\bar{C}(t,k)/\bar{C}(0,k)$ in the $(\log(t),\log(k))$ plane. Indeed, scaling in this representation simply corresponds to equal values of $\bar{C}(t,k)/\bar{C}(0,k)$ along straight lines.
Figure 4: Contour plots of the correlation function $\bar{C}(t,k)/\bar{C}(0,k)$ represented in the natural logarithm $(\log(k),\log(t))$ plane; (left): inviscid scaling obtained at $\nu=0$; (right): viscous scaling obtained at $\nu=0.016$. $\bar{C}(t,k)/\bar{C}(0,k)$ contour levels are drawn from $0.0$ to $0.9$ and spaced by $0.1$. The dashed black lines indicate (left) $t\sim k^{-1}$ and $t\sim k^{-2}$; (right) $t\sim k^{-2}$. The points $\tau_{1/2}(k)$, computed independently using $\bar{C}(\tau_{1/2},k)={\frac{1}{2}}\bar{C}(0,k)$ are indicated by red circles.
Figure 5: Contour plots of $\omega S(\omega,k)$, where $S(\omega,k)$ is the power spectrum, i.e the Fourier transform of $\bar{C}(t,k)/\bar{C}(0,k)$, represented in the natural logarithm $(\log(k),\log(\omega))$ plane; (left): inviscid scaling obtained at $\nu=0$; (right): viscous scaling obtained at $\nu=0.016$. The dashed black lines indicate (left) $\omega\sim k^{1}$ and $\omega\sim k^{2}$; (right) $\omega\sim k^{2}$. The contour levels of $\omega S(k,\omega)$ are: (left) from $0.0$ to $0.09$ (spaced by $0.01$) and (right) $0.0$ to $0.9$ (spaced by $0.1$).
The red circles  [^2] Red circles standing right on top of the $0.5$ contour line validates the interpolation scheme used to obtain the logarithmic representation from the raw numerical data that is equally spaced both in time and wavenumber. indicate the points $\tau_{1/2}(k)$ obtained independently by solving for each value of $k$ the implicit definition $\bar{C}(\tau_{1/2}(k),k)={\frac{1}{2}}\bar{C}(0,k)$. The left panel of Fig. 4 corresponds to the inviscid case $\nu=D=0$. The black dashed lines indicate the $t\sim k^{-1}$ and $t\sim k^{-2}$ scaling laws, which are both observable. In fact, the small-time behaviour of the correlation scaling with timescale $t\sim k^{-1}$ was already predicted in the inviscid limit in Ref. [12, 13], in agreement with the FRG exact result for small time delays (24). This scaling is apparent on the level lines from 0.9 to 0.5 which closely follow this law. Interestingly, at larger times, another scaling emerges, indicated by the level line $0.1$ which is definitely steeper and nearer to a $t\sim k^{-2}$ behaviour. This is an evidence for the FRG prediction for the large time delays behaviour (24). On the right panel of Fig. 4, the forced Navier-Stokes equation obtained at $\nu=0.016$ is displayed. The black dashed line indicates the theoretical scaling $t\sim k^{-2}$, corresponding to the diffusive FNS behaviour, which is well-followed for this large viscosity.
The scaling behaviour can also be observed on the spectrum $S(k,\omega)$, defined as the Fourier transform in time of $\bar{C}(t,k)/\bar{C}(0,k)$. The numerical noise on this quantity can be reduced by considering $\omega S(k,\omega)$ instead [57, 58]. Fig. 5 shows the contour plot of $\omega S(k,\omega)$. Scaling laws again correspond to straight lines in the $(\log(\omega),\log(k))$ plane. The left panel of Fig. 5 displays the inviscid ($\nu=0$) scalings, where the black dashed lines indicate $\omega\sim k$ and $\omega\sim k^{2}$ scaling laws. It is apparent that both scalings are present: $\omega\sim k^{2}$ for small $\omega$ and $\omega\sim k$ for large $\omega$. On the right panel, the viscous scaling obtained at $\nu=0.016$ is displayed. The black dashed line indicate the expected theoretical $\omega\sim k^{2}$ diffusive law, which is again well followed.
One can study as well the emergence of the $z=1$ scaling varying the viscosity. As highlighted in Fig. 3 obtained with FRG, one expects that at finite but non-zero viscosity, this scaling also appears at large wavenumbers, on a range all the more extended that the viscosity is small. We show on the left panel of Fig. 6 the half decay time $\tau_{1/2}(k)$ obtained from numerical simulations from $\nu=0.016$ to $\nu=0$, and on the right panel the same data compensated by $k$, i.e $k\tau_{1/2}(k)$, such that the $z=1$ scaling appears in this representation as a plateau. Note that at the highest momenta, all the curves saturate due to truncation effects. For the largest viscosities, only the diffusive regime $z=2$ is observed, while as the viscosity decreases, the behaviour of the large momentum mode bends towards a $z=1$ scaling. The plateau is visible on the right panel for viscosities $\nu\lesssim 0.0022$. The numerical results are thus in perfect accordance with the FRG prediction.
Figure 6: Left panel: half decay time $\tau_{1/2}$ of the correlations versus wavenumber $k$ for different viscosities obtained from numerical simulations. At the largest viscosity, only the viscous scaling $\tau_{1/2}\sim k^{-2}$ is visible, indicated by the solid line. At the smallest viscosities, the behaviour at large wavenumbers shows the inviscid scaling $\tau_{1/2}\sim k^{-1}$, indicated by the dashed line. Right panel: same as in left panel, but compensated by $k$, i.e $k\,\tau_{1/2}(k)$ versus $k$. Markers correspond on both panels to: $\nu=0.016$: $o$, $\nu=0.0044$: asterisk, $\nu=0.0022$: $+$, $\nu=0.0011$: diamond, $\nu=0.00055$: x and $\nu=0$: squares (colors from red to blue).
5 Conclusion
In this work, we have investigated the stochastic 3D Navier-Stokes equation in the presence of thermal noise, which corresponds to model A of FNS, using both direct numerical simulations and functional renormalisation group. We have identified a new fixed point, called IFNS, found in the inviscid (Euler) limit of model A, and which controls the large momentum behaviour of the correlation function. This fixed point is characterised by a critical dynamical exponent $z=1$ at small time delays. Moreover, the FRG result shows that the behaviour of the temporal correlations crossover at large time delays to a much slower decay with $z=2$. This second time regime had remained elusive in direct numerical simulations so far due to the difficulty to properly resolve long time delays. Our numerical simulations of model A have allowed us to observe both regimes, yielding the first numerical evidence of the long-time regime for 3D Navier-Stokes equation. The simulations have thus fully confirmed the FRG predictions.
Remarkably, the unveiled scaling $z=1$ is common to the inviscid regime of both the Burgers and Navier-Stokes equations. Moreover, the value of $z$ is independent of the dimension, as well as of the precise correlations of the stochastic term $\mathbf{f}$ provided it has a Gaussian distribution [34, 55]. Thus it holds both for equilibrium situations (when $\mathbf{f}$ is a thermal noise) and non-equilibrium situations (when $\mathbf{f}$ is a large-scale forcing), even though the corresponding static behaviours, eg. the kinetic energy spectra, are completely different. We have provided the explanation of this robustness, as we have shown that the large momentum behaviour of correlation functions is completely fixed and solely determined by the symmetries. Indeed, all these hydrodynamical equations share the same fundamental (extended) symmetries: the time-dependent Galilean symmetry and time-dependent shifts of the response fields (although the latter bears slightly different forms in Burgers and NS equations). These symmetries, through the related Ward identities, allow for an exact closure of the FRG flow equations in the asymptotic regime of large momenta. Their solution endows identical forms for Burgers or NS flows, differing only by non-universal constants, and featuring the exact $z=1$ exponent. The FRG has thus proved very efficient to understand the temporal behaviour of correlation functions. In perspective, it would be valuable to use it to gain more understanding of static properties of structure functions, and in particular intermittency effects.
Acknowledgements
LC wishes to thank G. Eyink for suggesting this work. LG acknowledges support by the MSCA Cofund QuanG (Grant Number : 101081458) funded by the European Union. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or Université Grenoble Alpes. Neither the European Union nor the granting authority can be held responsible for them.